OpenStudy (anonymous):

determine all values of c so that the lines cy+4x=9 and cy-5x=10 are perpendicular. Check solution

OpenStudy (cwrw238):

for 2 lines to be perpendicular their product of their slopes will be = -1

OpenStudy (cwrw238):

the slope of the first line is -4/c and of the second is 5/c so -4/c * 5/c = -1 for perpendicularity solve for c

OpenStudy (anonymous):

ok i got the answer im just not entirely sure how to check the solution

OpenStudy (cwrw238):

right first divide -4 by your answer then 5 by your answer to get the 2 slopes if correct then -1 divided by 1 slope should equal other slope

OpenStudy (cwrw238):

- got 2 values for c +/- sqrt20 or +/- 4.47214

OpenStudy (anonymous):

ok so the answer would be c=0 +/- 2 and square root of 5 how do i check if this is the correct solution?

OpenStudy (cwrw238):

right 2 sqrt5 = sqrt20 so one slope would be -4/2sqrt5 and the other is 5 / 2 sqrt5 -1/ first should work out to second use your calculator

OpenStudy (cwrw238):

- if the lines are perpendicular then -1 divided by one slope will equal other slope

OpenStudy (cwrw238):

i got -4 / (2sqrt5) = -0.894427191 5 / (2sqrt5) = 1.118033989 and -1 / 1.118033989 = -0.894427191

OpenStudy (anonymous):

oh okay thanks. I see how you did it

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

suppose a line goes through the point (3,-7) and is parallel to the line passing through the points (-1,8) and (1,2). Write equation of the line and determine the intercepts. Check solution

OpenStudy (cwrw238):

first find the slope of the line passing through the points (-1,8) and (1,2) then this will be the slope m going through the line you require then plug in value of m and the point (3,-7) into the standar form y - y1 = m(x - x1)